# Restoring Accessibility During Urban Rail Disruptions via Bus Network Redesign

### v1

@@ 5.2 Performance of Our Solution
we calculate this improvement, by assuming we add a certain number of extra buses (we test $0,5$ and $10$ extra buses). Note that, when the number of extra buses is 00, the Conventional Replacement Bus scenario corresponds exactly to the Before Disruption scenario (indeed, running a bus line that replaces the rail line with 00 buses operating on it is equivalent to operate no remediation at all). This suggests that our approach improves the accessibility of the vast majority of affected centroids, especially those that are heavily impacted (on the left of the $x$-axis). Even without adding any extra buses, the improvement of our solution over the classic approach goes up to 60% in the most affected centroids. Fig. 6(a) and Fig. 6(b) show the accessibility recovered by the Conventional Replacement Bus method and by our solution. In particular, let $\overline{acc}(\pazocal G)=\frac{1}{|\pazocal C|}\cdot\sum_{c_{i}\in\pazocal C}acc(c_{i},\pazocal G)$ be the average accessibility of PT configuration $\pazocal G$. We compute in Fig. 6 the following quantities: $\overline{acc}(\pazocal G^{\text{REPL}})/\overline{acc}(\pazocal G^{\text{O}})$ and $\overline{acc}(\pazocal G^{\text{OURS}})/\overline{acc}(\pazocal G^{\text{O}})$. For both Évry-Courcouronnes and Choisy-le-Roi cases, our method is consistently more efficient than the Conventional Replacement Bus method, even when we compare our solution with 0 extra buses with the Conventional Replacement Bus method with 10 extra buses.
@@
Fig. 7 shows the cumulative distribution functions (ECDF) of centroids’ accessibility across different scenarios. The green curves represent our method, and the red curves represent the conventional replacement bus strategy. We observed that our method substantially increases accessibility for all numbers of extra buses (10, 20, and 30 vehicles). For the conventional replacement buses, we can see that even if 30 additional buses are added, it is still not enough to recover the loss in accessibility. Our solution outperforms its results, even with 00 for additional buses. Moreover, adding extra buses in the conventional replacement bus start only brings negligible improvement, while our solution can consistently increase accessibility when increasing the fleet. Overall, our solution is more effective in narrowing the gap between before-disruption and after-remediation accessibility distribution.
Finally, we study the operational distance covered by the entire fleet of buses in the unit of time. As concerns the Conventional Replacement Bus strategy, we include in the calculation the distance covered by all the regular lines plus the distance covered by the replacement buses. As concerns our solution, instead, we include in the calculation the distance covered by all the regular lines plus the distance covered by the extended lines. We obtained real-world travel distances using Open-Street Map, assuming a set of lines $\pazocal{L}$, Total travel distance per unit of time = $\sum_{l\in\pazocal{L}}$ Circle length of line $l\times$ Number of service runs on line $l$ per unit of time. As shown in Fig. 8, our method significantly reduces the total operational distance in both the Évry-Courcouronnes and Choisy-le-Roi case studies. This suggests that our method reduces operational cost [40] and environmental impact.
@@ 6 Conclusion
This paper presents a method to redesign bus lines, calculating routes for bus line extensions and allocating buses among lines, to be applied in case of urban rail disruption. Its novelty is that, instead of considering conventional bus replacement approaches or bus bridging, we aim to restore the accessibility loss induced by the disruption.
@@ Acknowledgements
Conflict of Interest
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.

### v2

@@ 5.2 Performance of Our Solution
we calculate this improvement, by assuming we add a certain number of extra buses (we test $0,5$ and $10$ extra buses). Note that, when the number of extra buses is 0, the Conventional Replacement Bus scenario corresponds exactly to the Before Disruption scenario (indeed, running a bus line that replaces the rail line with 0 buses operating on it is equivalent to operate no remediation at all). This suggests that our approach improves the accessibility of the vast majority of affected centroids, especially those that are heavily impacted (on the left of the $x$-axis). Even without adding any extra buses, the improvement of our solution over the classic approach goes up to 60% in the most affected centroids. Figure 5 explicitly shows that compared to the traditional baseline (conventional replacement bus), our method reduces accessibility degradation impact for each centroid (with few exceptions), confirming that while travel time increases may occur, their negative impact on accessibility is mitigated. Fig. 6(a) and Fig. 6(b) show the accessibility recovered by the Conventional Replacement Bus method and by our solution. In particular, let $\overline{acc}(\pazocal G)=\frac{1}{|\pazocal C|}\cdot\sum_{c_{i}\in\pazocal C}acc(c_{i},\pazocal G)$ be the average accessibility of PT configuration $\pazocal G$. We compute in Fig. 6 the following quantities: $\overline{acc}(\pazocal G^{\text{REPL}})/\overline{acc}(\pazocal G^{\text{O}})$ and $\overline{acc}(\pazocal G^{\text{OURS}})/\overline{acc}(\pazocal G^{\text{O}})$. For both Évry-Courcouronnes and Choisy-le-Roi cases, our method is consistently more efficient than the Conventional Replacement Bus method, even when we compare our solution with 0 extra buses with the Conventional Replacement Bus method with 10 extra buses.
Fig. 7 shows the cumulative distribution functions (ECDF) of centroids’ accessibility across different scenarios. The green curves represent our method, and the red curves represent the conventional replacement bus strategy. We observed that our method substantially increases accessibility for all numbers of extra buses (10, 20, and 30 vehicles). For the conventional replacement buses, we can see that even if 30 additional buses are added, it is still not enough to recover the loss in accessibility. Our solution outperforms its results, even with 0 for additional buses. Moreover, adding extra buses in the conventional replacement bus start only brings negligible improvement, while our solution can consistently increase accessibility when increasing the fleet. Overall, our solution is more effective in narrowing the gap between before-disruption and after-remediation accessibility distribution.
Finally, we study the operational distance covered by the entire fleet of buses in the unit of time. As concerns the Conventional Replacement Bus strategy, we include in the calculation the distance covered by all the regular lines plus the distance covered by the replacement buses. As concerns our solution, instead, we include in the calculation the distance covered by all the regular lines plus the distance covered by the extended lines. We obtained real-world travel distances using Open-Street Map, assuming a set of lines $\pazocal{L}$, Total travel distance per unit of time = $\sum_{l\in\pazocal{L}}$ Circle length of line $l\times$ Number of service runs on line $\ell$ per unit of time. As shown in Fig. 8, our method significantly reduces the total operational distance in both the Évry-Courcouronnes and Choisy-le-Roi case studies. This suggests that our method reduces operational cost [40] and environmental impact.
While our optimized bus configuration demonstrates superior accessibility restoration compared to conventional methods, it serves as a temporary adaptation strategy. The rail service should ultimately be restored due to its inherent advantages in passenger capacity, dedicated rights-of-way, and environmental benefits. However, the accessibility insights gained could inform future network resilience planning, particularly for areas with frequent maintenance or recurring disruptions.
@@ 6 Conclusion
This paper presents a method to redesign bus networks during urban rail disruptions by calculating optimal routes for bus line extensions and allocating vehicles among lines. Unlike conventional bus replacement or bridging approaches that focus on operational metrics, our method specifically aims to restore accessibility loss induced by the disruption.